OpenStudy (anonymous):

. Find the slope of the line containing the given pair of points. (14.4, -7.8) and ( -12.5, -17.6)

OpenStudy (anonymous):

Slope equation \[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (katrinakaif):

9.8/26.9

OpenStudy (callisto):

slope=[(-7.8)-(-17.6)]/[(14.4 - (-12.5)] = 9.8/26.9